OpenStudy (anonymous):

what is the LCM of 7(t-2) and 14(t-2) what is the LCM of 7(t-2) and 14(t-2) @Mathematics

myininaya (myininaya):

14(t-2) since 7(t-2) times 2=14(t-2) and 14(t-2) times 1=14(t-2)

OpenStudy (anonymous):

so would it be (t-2) as the lcm?

myininaya (myininaya):

no 14(t-2)

OpenStudy (anonymous):

ok thanks